# Privacy vs. Profit: The Impact of Google’s Manifest Version 3 (MV3) Update on Ad Blocker Effectiveness

### v1

@@ Abstract
Google’s recent update to the manifest file for Chrome browser extensions—transitioning from manifest version 2 (MV2) to manifest version 3 (MV3)—has raised concerns among users and ad blocker providers, who worry that the new restrictions, notably the shift from the powerful WebRequest API to the more restrictive DeclarativeNetRequest API, might reduce ad blocker effectiveness. Because ad blockers play a vital role for millions of users seeking a more private and ad-free browsing experience, this study empirically investigates how the MV3 update affects their ability to block ads and trackers.
Through a browser-based experiment conducted across multiple samples of ad-supported websites, we compare the MV3 to MV2 instances of four widely used ad blockers. Our results reveal no statistically significant reduction in ad-blocking or anti-tracking effectiveness for MV3 ad blockers compared to their MV2 counterparts, and in some cases, MV3 instances even exhibit slight improvements in blocking trackers. These findings are reassuring for users, indicating that the MV3 instances of popular ad blockers continue to provide effective protection against intrusive ads and privacy-infringing trackers. While some uncertainties remain, ad blocker providers appear to have successfully navigated the MV3 update, finding solutions that maintain the core functionality of their ad blockers.
@@ 1 Introduction
The MV3 update restricts the ability of ad blockers to intercept and modify network requests, which has raised concerns about their possibly reduced effectiveness in blocking ads and trackers. Nevertheless, since the MV3 update has been announced, ad blocker providers have worked intensely to close the gap and provide a seemingly identical experience under the new restrictions. Despite these efforts users remain concerned about the impact of the MV3 update on ad blocking effectiveness and consequently their privacy [11, 12].
These user concerns show that Google’s underlying motivation for the update remains opaque, primarily because its internal decision-making processes and strategic incentives are not fully visible to the public. The MV3 update might deliberately weaken the ad-blocking and anti-tracking effectiveness of ad blockers, thereby consolidating Google’s control over user data and increasing its ad revenue. This outcome would align with patterns in digital advertising, where companies actively work to bypass user interventions—like ad blockers—to protect and expand their monetization potential [13].
Alternatively, the MV3 update could represent a strategic move toward increased user privacy. By reducing potential avenues for intrusive data collection, Google might aim to strengthen user trust—a factor that research in privacy economics (e.g., [13]) suggests can lead to more user engagement. Such research has shown that enhanced privacy protections are often correlated with increased platform loyalty, as users tend to favor services that safeguard their data.
@@
The second research question builds on this comparison by examining the effectiveness of individual MV3 ad blockers compared to their MV2 counterparts. Since most ad blocker providers state that their MV3 instances use either a reduced or modified set of blocking rules derived from their MV2 instances [14, 15, 16, 17], MV2 instances of ad blockers serve as natural benchmarks. By directly comparing each MV3 ad blocker to its MV2 counterpart—including Adblock Plus, AdGuard, Stands, and uBlock Origin—we assess how ad blocker providers have adapted to the MV3 update. This comparison is particularly relevant for users transitioning to MV3, as it provides insights into whether some MV3 ad blockers remain as effective as their MV2 counterparts, while others may exhibit reduced effectiveness. Thus, the second research question aims to help users choose the most effective MV3 ad blocker amid the MV2 to MV3 update. It also evaluates whether different ad blocker providers have adjusted differently to the MV3 update [16, 18, 19, 20, 21].
The third research question addresses whether using multiple MV3 ad blockers simultaneously impacts ad-blocking and anti-tracking effectiveness compared to using individual MV3 ad blockers. This research question is relevant because it addresses concerns expressed by ad blocker providers about the potentially reduced effectiveness of the MV3 ad blocker when it is used simultaneously with multiple other MV3 ad blockers [22]. The third research question also measures the effectiveness of MV3 ad blockers for a privacy-aware user who uses multiple MV3 ad blockers simultaneously.
Thus, this study explores the impact of Google’s update from MV2 to MV3 on the effectiveness of ad blockers in blocking ads and trackers. Through a browser-based experiment, we compare MV3 ad blockers to their MV2 counterparts and examine the effect of using multiple MV3 ad blockers simultaneously. We initiate the browser-based experiment on a main sample of 114 websites across ten separate measurement runs, and two additional samples of 191 websites and 185 websites stratified by website employee count and popularity rank, respectively. In addition, we run the browser-based experiment using Chrome and Firefox on our main sample of 114 websites, to ensure the cross-browser robustness of our results. We complement these experiments with additional analyses including alternative effectiveness metrics, assessments of early MV3 implementations, and detailed visual inspections of website screenshots—that reinforce the robustness and generalizability of our findings.
@@
  • Empirical comparison between MV3 and MV2 groups of ad blockers: This study presents the first empirical analysis to evaluate the effectiveness of MV3 ad blockers compared to their MV2 counterparts. The results reveal no statistically significant reduction in ad-blocking or anti-tracking effectiveness between the MV3 and the MV2 groups of ad blockers.
  • Variability among individual MV3 ad blockers: When comparing individual ad blockers, we find no significant differences between MV3 and MV2 instances for all ad blockers—except for the Stands ad blocker. The Stands MV3 ad blocker blocks 6 more trackers than its MV2 counterpart, indicating a 30.5% improvement in anti-tracking effectiveness compared to its MV2 counterpart.
  • Enhanced anti-tracking effectiveness with multiple MV3 ad blockers: Using multiple MV3 ad blockers simultaneously does not impact their ad-blocking effectiveness. However, it significantly enhances anti-tracking effectiveness compared to using some MV3 ad blockers alone. Specifically, the combination blocks around 13 more trackers (61.8%) than Adblock Plus MV3 alone, 7 more trackers (25.4%) than AdGuard MV3 alone, and 7 more trackers (26.9%) than Stands MV3 alone. This improvement in anti-tracking effectiveness is driven by the inclusion of uBlock MV3 in the combination, suggesting a strategic advantage in combining certain MV3 ad blockers to improve users’ online privacy. This result alleviates concerns of ad blocker providers that using multiple MV3 ad blockers simultaneously reduces their effectiveness.
@@ 2.2 Manifest Version 2 (MV2) Ad Blocker
Once a user installs an ad blocker, the user grants it permission to monitor and block HTTP requests. In simple terms, the ad blocker contains a list of blocking rules. As the user’s browser makes each HTTP request, the ad blocker checks it against that list of blocking rules and blocks the request if it finds a match. Under Google’s MV2, ad blockers use the WebRequest API and background pages to monitor real-time HTTP requests between a user’s browser and the website they visit.
@@ 3.1 Description of Google’s MV3 Update From Two Opposing Perspectives
Google announced the MV3 update in November 2018 and made it available to extension developers in beta form in December 2020 [7, 9]. In January 2022, Google stopped allowing developers to add new MV2 extensions to the CWS, signaling the start of the MV2 phase out [28]. In late 2024, Google warned developers that it intended to complete the MV2 phase out by June 2025 [29], although the company may still delay the final phase out.
@@
By introducing the MV3 update, Google states that it seeks to reduce the power of extensions and enhance Chrome browser performance, thereby restricting the potential misuse of user data by malicious extensions. The two critical technical changes in the MV3 update reflect Google’s effort to enhance security and privacy of extensions [32]. The first change replaces the WebRequest API, which allows ad blockers to monitor and block ads and trackers in real-time, with the DeclarativeNetRequest API. This change restricts ad blockers to operating based on a predefined set of up to 30,000 blocking rules, while allowing additional flexibility through updated blocking rule limits. The second change, transition from background pages to service workers, intends to improve extension performance by managing background tasks more efficiently.
@@
Further, replacing background pages with service workers to improve browser performance raises practical concerns for ad blocker providers. They argue that service workers can disrupt the continuous monitoring of HTTP requests, which might lead to "ad flickering" (i.e., short re-appearances of website ads) and put ad blockers "[…] into a kind of sleep mode" [22]. Ad blocker providers also claim that their ad blockers are already highly performant and do not significantly impact browser speed, questioning the value of Google’s change in terms of improving browser performance [35]. Borgolte and Feamster [23] addressed that concern and empirically demonstrated that MV2 ad blockers do not significantly impact browser speed.
Thus, while Google’s MV3 update seeks to enhance user privacy, security, and potentially Chrome browser performance, it introduces changes that raise concerns for ad blocker providers. The new restrictions and technical requirements challenge their ability to block ads and trackers effectively. As the deadline for Google’s phase out of MV2 extensions approaches, ad blocker providers are grappling with innovating and adapting to changes introduced with MV3’s update to maintain their role in providing users with improved privacy and ad-free browsing experiences [16, 18, 21, 36, 20].
@@ 3.2 Changes and Workarounds to Google’s MV3 Update
Despite these changes and workarounds, ad blocker providers continue to warn about potentially reduced core functionality [14, 15, 16, 17]. Whether these adaptations have fully alleviated the technical challenges introduced by the MV3 update remains uncertain. The MV3 update could lead to reduced ad blocker effectiveness, no actual impact, or even improvements—though the latter would be unexpected.
@@
The unexpected scenario of a positive impact—where ad blocker providers adapt to MV3’s restrictions and even innovate to improve effectiveness—could lead to more effective ad blocking, enhancing user privacy and creating an even more ad-free browsing experience. Users would encounter fewer ads and fewer trackers. Ad blocker providers would not face increased development costs while demonstrating strong adaptability and innovation. In contrast, Google and website publishers would experience a decline in ad impressions, potentially sacrificing some ad revenue while reinforcing user trust.
@@ 4 Related Work
Garimella et al. [24] expanded on Wills and Uzunoglu [26], highlighting uBlock’s superior ad-blocking and anti-tracking effectiveness, as compared with other ad blockers. Their study showed a paradox of using ad blockers for user privacy: some can even add extra trackers on websites users visit, thereby harming it instead of improving it.
@@ 5.1 Selection of MV3 and MV2 Ad Blockers
We used the current versions of each ad blocker available as of August 16, 2024. According to CWS statistics, more than 133 million users actively use these eight ad blockers worldwide, with significantly fewer users currently using their MV3 instances—which reflects the study’s relevance to a substantial user base. Moreover, the anticipation of a user base transition from MV2 to MV3 instances, as Google plans to fully phase out MV2 extensions by June 2025, further emphasizes the timeliness of this study in evaluating ad blocker effectiveness.
@@ 5.2 Description of Browser Experiment and Data Collection Process
We conducted a browser-based experiment to translate that scenario into a controlled experimental setup (Figure 2). To this end, we used the Google Chrome browser, reflecting the everyday user environment on a local machine with a European IP address. The choice of a local machine over a cloud platform (e.g., Google Cloud Platform) mirrors the typical user. Past research (e.g, [23]) indicates no significant difference between a local machine and a cloud platform in such a web measurement. We used a European IP address, which provides a conservative estimate of ad blocker effectiveness. Further, we automated the experiment using Selenium, a Python-based package that uses a user-agent based on the Chrome browser but also includes an indication of automation. This user-agent ensures that, in most cases, the webpages serve content as they would to a regular user.
@@
We ran ten separate browser instances in parallel, each one representing a unique user—one without an ad blocker and the others with a selection of ad blockers from Adblock Plus, AdGuard, Stands, and uBlock, including instances of MV2, MV3, and a combination of MV3 ad blockers (MV3+). This parallel setup was crucial to maintaining uniform testing conditions and to measure each ad blocker’s ad-blocking and anti-tracking effectiveness under identical web browsing conditions. Running the instances in parallel ensures that all ad blockers are tested under the same conditions simultaneously.
@@
We launched the browser-based experiment on November 27, 2024. To account for webpage dynamism, we conducted ten separate measurement runs. A descriptive analysis revealed little variance between those runs (see Appendix A.2), and we averaged the results for our analysis. Each website crawl took about 2 minutes on average. The total duration of the experiment for our main sample was around 60 hours and 54 minutes.
To ensure the cross-browser applicability of our findings, we also launched the browser-based experiment using the Firefox browser. However, conducting the browser-based experiment using the Firefox browser presents a key challenge: unlike in Chrome browser, we could not use the same set of MV3 and MV2 ad blockers. The Firefox browser has announced that it will continue supporting both MV3 and MV2 extensions in the future [46] and currently has no plans to phase out MV2 extensions. While this strategy ensures ongoing availability of MV2 ad blockers in the Firefox browser, it has also delayed the full adaption of ad blocker providers to MV3. As a result, only uBlock Origin provides both MV3 and MV2 instances in the Firefox browser. Complicating matters further, as of late 2024, the Firefox Add-ons Store [^4] https://addons.mozilla.org/en-US/firefox/ removed uBlock Origin Lite—uBlock’s MV3 instance—due to disagreements with its developer [47]. Although no longer listed in the Firefox Add-ons Store, uBlock Origin Lite remains available for manual installation in the Firefox browser. Despite these challenges, we conducted a limited version of our browser-based experiment using the Firefox browser to ensure the cross-browser applicability of our findings (see Section 6.4 and Appendix B.4).

### v2

@@ Abstract
Google’s recent update to the manifest file for Chrome browser extensions, transitioning from manifest version 2 (MV2) to manifest version 3 (MV3), has raised concerns among users and ad blocker providers, who worry that the new restrictions, notably the shift from the powerful WebRequest API to the more restrictive DeclarativeNetRequest API, might reduce ad blocker effectiveness. Because ad blockers play a vital role for millions of users seeking a more private and ad-free browsing experience, this study empirically investigates how the MV3 update affects their ability to block ads and trackers.
Through a browser-based experiment conducted across multiple samples of ad-supported websites, we compare the MV3 to MV2 instances of four widely used ad blockers. Our results reveal no statistically significant reduction in ad-blocking or anti-tracking effectiveness for MV3 ad blockers compared to their MV2 counterparts, and in some cases, MV3 instances even exhibit slight improvements in blocking trackers. These findings are reassuring for users, indicating that the MV3 instances of popular ad blockers continue to provide effective protection against intrusive ads and privacy-infringing trackers. While some uncertainties remain, ad blocker providers appear to have successfully navigated the MV3 update, finding solutions that maintain the core functionality of their extensions.
Note. This is the camera-ready version accepted to Proceedings on Privacy Enhancing Technologies (PoPETs), Volume 2026, Issue 1. The official version with page numbers and DOI will appear on the PoPETs website when the 2026(1) issue is released.
Please cite as:
K. Lukic and L. Papadopoulos. “Privacy vs. Profit: Impact of Google’s MV3 on Ad-Blocker Effectiveness.” Proceedings on Privacy Enhancing Technologies, 2026(1) (forthcoming).
@@ 1 Introduction
The MV3 update restricts the ability of ad blockers to intercept and modify network requests, which has raised concerns about their possibly reduced effectiveness in blocking ads and trackers. Nevertheless, since the MV3 update has been announced, ad blocker providers have worked intensely to close the gap and provide a seemingly identical experience under the new restrictions. Despite these efforts users remain concerned about the impact of the MV3 update on ad-blocking effectiveness and consequently their privacy [11, 12].
These user concerns show that Google’s underlying motivation for the update remains opaque, primarily because its internal decision-making processes and strategic incentives are not fully visible to the public. The MV3 update might deliberately weaken the ad-blocking and anti-tracking effectiveness of ad blockers, thereby consolidating Google’s control over user data and increasing its ad revenue. This outcome would align with patterns in digital advertising, where companies actively work to bypass user interventions, like ad blockers, to protect and expand their monetization potential (e.g., [13]).
Alternatively, the MV3 update could represent a strategic move toward increased user privacy. Previous research has shown that only a small number of blocking rules are effectively used by ad blockers [14], leaving much room for optimization. Further, by reducing potential avenues for intrusive data collection, Google might aim to strengthen user trust, a factor that research in privacy economics (e.g., [13]) suggests can lead to more user engagement. Such research has shown that enhanced privacy protections are often correlated with increased platform loyalty, as users tend to favor services that safeguard their data.
@@
The second research question builds on this comparison by examining the effectiveness of individual MV3 ad blockers compared to their MV2 counterparts. Since most ad blocker providers state that their MV3 instances use either a reduced or modified set of blocking rules derived from their MV2 instances [15, 16, 17, 18], MV2 instances of ad blockers serve as natural benchmarks. By directly comparing each MV3 ad blocker to its MV2 counterpart, including Adblock Plus, AdGuard, Stands, and uBlock Origin, we assess how ad blocker providers have adapted to the MV3 update. This comparison is particularly relevant for users transitioning to MV3, as it provides insights into whether some MV3 ad blockers remain as effective as their MV2 counterparts, while others may exhibit reduced effectiveness. Thus, the second research question aims to help users choose the most effective MV3 ad blocker amid the MV2 to MV3 update. It also evaluates whether different ad blocker providers have adjusted differently to the MV3 update [17, 19, 20, 21, 22].
The third research question addresses whether using multiple MV3 ad blockers simultaneously impacts ad-blocking and anti-tracking effectiveness compared to using individual MV3 ad blockers. This research question is relevant because it addresses concerns expressed by ad blocker providers about the potentially reduced effectiveness of the MV3 ad blocker when it is used simultaneously with multiple other MV3 ad blockers [23]. This concern is substantiated by the global rule limit introduced with the MV3 update. The third research question also measures the effectiveness of MV3 ad blockers for a privacy-aware user who uses multiple MV3 ad blockers simultaneously.
Thus, this study explores the impact of Google’s update from MV2 to MV3 on the effectiveness of ad blockers in blocking ads and trackers. Through a browser-based experiment, we compare MV3 ad blockers to their MV2 counterparts and examine the effect of using multiple MV3 ad blockers simultaneously. We initiate the browser-based experiment on a main sample of 924 websites across five separate measurement runs, and two additional samples of 191 websites and 185 websites stratified by website employee count and popularity rank, respectively. In addition, we run the browser-based experiment using Chrome and Firefox on our main sample of 924 websites, to ensure the cross-browser robustness of our results. We complement these experiments with additional analyses including alternative effectiveness metrics, assessments of early MV3 implementations, and detailed visual inspections of website screenshots that reinforce the robustness and generalizability of our findings.
@@
  • Empirical comparison between MV3 and MV2 groups of ad blockers: This study presents the first empirical analysis to evaluate the effectiveness of MV3 ad blockers compared to their MV2 counterparts. The results reveal no statistically significant reduction in ad-blocking effectiveness and a statistically significant increase in anti-tracking effectiveness, with the MV3 group blocking about 1.8 more trackers per website on average than the MV2 group of ad blockers.
  • Variability among individual MV3 ad blockers: When comparing individual ad blockers, we find no significant differences between MV3 and MV2 instances for AdGuard and uBlock; however, Adblock Plus MV3 blocks about 1.9 more trackers than its MV2 counterpart (21.5%), and Stands MV3 blocks about 5.2 more trackers (45.9%) than its MV2 counterpart.
  • Enhanced anti-tracking effectiveness with multiple MV3 ad blockers: Using multiple MV3 ad blockers simultaneously does not impact their ad-blocking effectiveness. However, it significantly enhances anti-tracking effectiveness compared to using some MV3 ad blockers alone. Specifically, the combination blocks about 10.3 more trackers (95.3%) than Adblock Plus MV3 alone, 4.4 more (26.6%) than AdGuard MV3 alone, and 4.6 more (27.8%) than Stands MV3 alone, with no difference relative to uBlock MV3. This improvement in anti-tracking effectiveness is driven by the inclusion of uBlock MV3 in the combination, suggesting a strategic advantage in combining certain MV3 ad blockers to improve users’ online privacy. This result alleviates concerns of ad blocker providers that using multiple MV3 ad blockers simultaneously reduces their effectiveness.
@@ 2.2 Manifest Version 2 (MV2) Ad Blocker
In simple terms, the ad blocker contains a list of blocking rules. As the user’s browser makes each HTTP request, the ad blocker checks it against that list of blocking rules and blocks the request if it finds a match. Under Google’s MV2, ad blockers use the WebRequest API and background pages to monitor real-time HTTP requests between a user’s browser and the website they visit.
@@ 3.1 Description of Google’s MV3 Update From Two Opposing Perspectives
Google announced the MV3 update in November 2018 and made it available to extension developers in beta form in December 2020 [7, 9]. In January 2022, Google stopped allowing developers to add new MV2 extensions to the CWS, signaling the start of the MV2 phase out [29]. In late 2024, Google warned developers that it intended to complete the MV2 phase out by July 2025 [30].
@@
By introducing the MV3 update, Google states that it seeks to reduce the power of extensions and enhance Chrome browser performance, thereby restricting the potential misuse of user data by malicious extensions. The two critical technical changes in the MV3 update reflect Google’s stated effort to enhance the security and privacy of extensions [33]. The first change replaces the WebRequest API, which allows ad blockers to monitor and block ads and trackers in real-time, with the DeclarativeNetRequest API. This change restricts ad blockers to operating based on a predefined set of up to 30,000 blocking rules. That is less than half the number of rules in EasyList [^1] https://easylist.to, the base filter set used by many extensions (e.g., uBlock Origin), and significantly restricts their processing pipeline. However, empirical analyses indicate that this constraint might not necessarily diminish ad blocker effectiveness [14, 34].
The second change, the transition from background pages to service workers, intends to improve extension performance by managing background tasks more efficiently. A service worker, unlike a more persistent background page, spins up only when an extension needs it (i.e., it is event-driven) and is torn down after 30 seconds of idle time. Because the service worker runs off the main thread and does not linger in memory, Chrome can reclaim RAM and CPU that dormant extensions once consumed, giving users a leaner, faster browser. The shorter life cycle might also reduce the window in which malicious or buggy code could abuse persistent privileges, so the change aims to strengthen security by design [35].
@@
Further, replacing background pages with service workers to improve browser performance raises practical concerns for ad blocker providers. They argue that short-lived service workers can disrupt the continuous monitoring of HTTP requests, which might lead to "ad flickering" (i.e., short re-appearances of website ads) and put ad blockers "[…] into a kind of sleep mode" [23]. More precisely, because Chrome automatically suspends an extension’s service worker after some time, all global variables, registered listeners and timers are wiped. That means an ad blocker must reload its state from storage and attach listeners immediately on wake-up or it will miss the very event that triggered the restart.
This loss of state and potential event-drop can interrupt or delay filtering, so ad blocker providers must redesign their code to persist these challenges [39]. Ad blocker providers also claim that their ad blockers are already highly performant and do not significantly impact browser speed, questioning the value of Google’s change in terms of improving browser performance [38]. Borgolte and Feamster [24] addressed that concern and empirically demonstrated that MV2 ad blockers do not significantly impact browser speed.
Thus, while Google’s MV3 update seeks to enhance user privacy, security, and potentially Chrome browser performance, it introduces changes that raise concerns for ad blocker providers. These new restrictions and technical requirements challenge their ability to block ads and trackers effectively. As Google phases out MV2 extensions, ad blocker providers are grappling with innovating and adapting to changes introduced with MV3’s update to maintain their role in providing users with improved privacy and ad-free browsing experiences [17, 19, 22, 40, 21].
@@ 3.2 Changes and Workarounds to Google’s MV3 Update
Despite these changes and workarounds, ad blocker providers continue to warn about potentially reduced core functionality [15, 16, 17, 18]. According to them, some problems persist. For example, the inability to define rules based on top-level context [16], a persistent limit of simultaneously active filter lists [17], or the fact that service workers might still be put to sleep, despite workarounds [15]. Thus, whether these adaptations have fully alleviated the technical challenges introduced by the MV3 update remains uncertain. The MV3 update could lead to reduced ad blocker effectiveness, no actual impact, or even improvements, though the latter would be unexpected.
@@
The unexpected scenario of a positive impact, where ad blocker providers adapt to MV3’s restrictions and even innovate to improve effectiveness, could lead to more effective ad blocking, enhancing user privacy and creating an even more ad-free browsing experience. Users would encounter fewer ads and fewer trackers. Ad blocker providers would not face increased development costs while demonstrating strong adaptability and innovation. In contrast, Google and website publishers would experience a decline in ad impressions, potentially sacrificing some ad revenue while reinforcing user trust.
3.3 Reactions of Other Browser Vendors
Many other browser vendors have approached the MV3 update with caution. Several leading ones such as Brave, Opera, Vivaldi, Firefox, and Safari accepted MV3 extensions while still allowing at least some extensions to run under MV2. For instance, Brave migrated to MV3 yet continues to expose AdGuard, uBlock Origin, uMatrix, and NoScript extensions through its own backend, reflecting a more supportive stance than in its initial roadmap [42]. Vivaldi, by contrast, states it is effectively obliged to phase out MV2 support, but emphasizes that its built-in anti-tracking and ad-blocking features should offset any major concerns regarding the MV3 update [43].
Safari [44] and Firefox [45], meanwhile, promise full ongoing support for both MV2 and MV3 extensions, and have announced no retirement date for MV2. On the other end of the spectrum, Microsoft Edge has been least skeptical from the outset: since July 2022 it has blocked new MV2-based submissions of extensions, arguing that this reduces fragmentation for developers [46], mirroring Google’s argumentation for the MV3 update.
In summary, because Chrome commands the dominant market share, many other browser vendors appear to have aligned themselves with Google’s new update, willingly or unwillingly.
@@ 4 Related Work
Garimella et al. [25] expanded on Wills and Uzunoglu [27], highlighting uBlock’s superior ad-blocking and anti-tracking effectiveness, as compared to other ad blockers. Their study showed a paradox of using ad blockers for user privacy: some can even add extra trackers on websites users visit, thereby harming it instead of improving it.
@@ 5.1 Selection of MV3 and MV2 Ad Blockers
We used the current versions of each ad blocker available as of August 16, 2024. According to CWS statistics, more than 133 million users actively use these eight ad blockers worldwide, with significantly fewer users currently using their MV3 instances—which reflects the study’s relevance to a substantial user base. Moreover, the anticipation of a user base transition from MV2 to MV3 instances, as Google plans to fully phase out MV2 extensions by July 2025, further emphasizes the timeliness of this study in evaluating ad blocker effectiveness.
@@ 5.2 Description of Browser Experiment and Data Collection Process
We conducted a browser-based experiment to translate that scenario into a controlled experimental setup (Figure 2). We ran all crawls on a local machine using the Google Chrome browser, reflecting everyday user practice, with a European residential IP address. Using a local machine kept the setup simple, avoided VPN or proxy artifacts, and held geography constant so that any differences reflect MV3 vs. MV2 rather than location. Because all measurements originated from a European residential IP, and GDPR typically reduces third-party activity (e.g., [54, 55, 56]) due to compliance risks irrespective of cookie consent, EU-origin visits tend to activate fewer third parties. As a result, the absolute ad and tracker counts we report are likely conservative relative to a non-EU vantage point. We automated the experiment using Selenium (Python) and a Chrome user-agent so that webpages would, in most cases, serve content as they would for a standard visitor.
@@
We ran ten separate browser instances in parallel, each one representing a unique user—one without an ad blocker and the others with a selection of ad blockers from Adblock Plus, AdGuard, Stands, and uBlock, including instances of MV2, MV3, and a combination of these MV3 ad blockers (MV3+). This parallel setup was crucial to maintaining uniform testing conditions and to measure each ad blocker’s ad-blocking and anti-tracking effectiveness under identical web browsing conditions. Running the instances in parallel ensures that all ad blockers are tested under the same conditions simultaneously.
@@
We launched the browser-based experiment on May 15, 2025. To account for webpage dynamism (e.g., [57]), we conducted five separate measurement runs (runs). Unless stated otherwise, all main sample analyses use per-website averages computed across these five runs for each browser instance; all statistical tests and error bars are based on these per-website averages. A descriptive analysis revealed low variance between runs (see Appendix A.2). Each run took about 64 hours on average. The total duration of the experiment for our main sample was around 320 hours and 11 minutes, driven by the extensive wait times and webpage scrolling needed for the consistent measurements.
To ensure the cross-browser applicability of our findings, we also launched the browser-based experiment using the Firefox browser. However, conducting the browser-based experiment using the Firefox browser presents a key challenge: unlike in Chrome browser, we could not use the same set of MV3 and MV2 ad blockers. Firefox has announced that it will continue supporting both MV3 and MV2 extensions in the future [58] and currently has no plans to phase out MV2 extensions. While this strategy ensures ongoing availability of MV2 ad blockers in the Firefox browser, it has also delayed the full adaption of ad blocker providers to MV3. As a result, only uBlock Origin provides both MV3 and MV2 instances in the Firefox browser. Complicating matters further, as of late 2024, the Firefox Add-ons Store [^5] https://addons.mozilla.org/en-US/firefox/ removed uBlock Origin Lite—uBlock’s MV3 instance—due to disagreements with its developer [59]. Although no longer listed in the Firefox Add-ons Store, uBlock Origin Lite remains available for manual installation in the Firefox browser. Despite these challenges, we conducted a limited version of our browser-based experiment using the Firefox browser to ensure the cross-browser applicability of our findings (see Section 6.4 and Appendix B.4). In this Firefox browser-based experiment, we evaluated uBlock MV3 and MV2 on a main sample of 1,000 websites across five measurement runs beginning July 12, 2025, and obtained consistent results for 824 websites (i.e., websites with valid data across all five runs; others were excluded due to redirects, anti-automation/ad-blocking, or failed instrumentation).